Fichas
Stats 151 - Mean, Standard Deviation, and Empire Rule for Random Variables and Binomial Distribution
Fichas del conjunto (16)
- 1how do you find the mean of a random variable?sum of all x values times the probability = Ux = E(x)*p(x)
- 2a gambler bets 10 dollars on red on a roulette wheel that have 18 red, 18 black, and 2 green slots. Whats his expected return?10*(18/38) + (-10)*(20/38) = -0.526
on average the gambler will loose 52.6 cents per bet - 3how do you find the standard deviation of a random variable?sigma = square root(E(x-u)squared * p(x))
- 4a gambler bets 10 dollars on a specific roulette number. he gets 350 if he wins with an average loss of 52.6 cents a bet. What is the standard deviation?sigma y = square root((350-(0.53)squared)*(1/38)+(-10 - (0.53))squared * 37/38 = 57.63
- 5what is the probability you will fall into the middle section of a bell curve?68% = 0.68 probability
- 6What is a Bernoulli experiment?any experiment that has only two outcomes e.g. yes or no
- 7how do you find the probability of the outcomes for a Bernoulli experimentp(outcome 1) = p
p(outcome 2) = 1-p - 8what is the mean of a bernoulli experiment? the sigma?the mean is p. sigma is found by calculating (p(1-p))
- 9what is a binomial experiment?a bernoulli experiment performed more then once
- 10what do x and n-x equal in a binomial experiment?x = # of success'
n-x = # of failures - 11How do you find the probability of x in a binomial experiment?(n over x) P to the x (1-p) to the n-x
- 12What are the Bernoulli random variables represented as? How do they appear on a number line?the variables are presented as E1 and E2. E1 becomes 0 on the number line and E2 becomes 1
- 13when you see n over x in brackets, what does it mean?nCx
- 14Suppose that revenue Canada reports that 8% of all tax returns contain errors. An experiment consists of checking errors in a random sample of 10 tax returns. Find the probability that the sample contains exactly one errorP(x=1) = (10 over 1)*0.8^(1)*(1-0.8)^(10-1) = 0.3777
- 15at most one error?P(x< or equal to 1) = (x=0) + P(X=1) = (10 over 0) * 0.8^(0)(1-0.8)^(10-0) +0.3777 = 0.4344 + 0.3777 = 0.8121
- 16at least one error?P(x>or equal to 1) = 1-P(x=0) = 1-0.4344 = 0.5656
Información
- Descripción
- Chemistry 101 Stats 151 Fichas sobre Stats 151 - Mean, Standard Deviation, and Empire Rule for Random Variables and Binomial Distribution, creado por jennabarnes12387 el 28/01/2014.
- Creado por
- jennabarnes12387
- Visitas
- 231
- Etiquetas
- Valoración